# Weyl law for 1-cycles

### v1

@@ 4.2. Two important propositions to go from disks to manifolds
Given $C\in\operatorname{Faces}_{p}(X)$, we choose $R_{C}\in[L+K\delta,L+(K+1)\delta]$ ($K=K(n)$ to be determined later) such that $\partial B_{R_{C}}(e)\cap B_{i}^{C}=\emptyset$ for every $i\in I_{C}$ and $\partial B_{R_{C}}(e)\cap\mathcal{S}=\emptyset$ (being $\{B_{i}^{C}\}_{i\in I_{C}}$ the $\delta$-admissible family associated to $C$ and $\mathcal{S}=\bigcup_{x\in X_{0}}\operatorname{spt}(F^{\prime}(x))$). We denote
$J_{C}=\{i\in I_{C}:B_{i}\cap B_{R_{C}}(e)=\emptyset\}.$
We claim that there exists a hyperplane $\Pi\subseteq\mathbb{R}^{n-1}$ such that $S_{\Pi}\cap B_{i}^{C}=\emptyset$ for every $i\in J_{C}$ and also $S_{\Pi}\cap\mathcal{S}=\emptyset$. Consider the map $\operatorname{proj}:\mathbb{S}^{n-1}_{R_{C}}\to\mathbb{S}^{n-2}_{R_{C}}$ given in $\phi$-coordinates by $\operatorname{proj}(u,t)=(u,-\cos(R_{C}))$. Notice that $\operatorname{proj}$ is $1$-Lipschitz and that $B_{i}^{C}\cap\mathbb{R}^{n}_{-}\subseteq\mathbb{S}^{n-1}_{R_{C}}$ for every $i\in J_{C}$. Therefore if $\tilde{B}_{i}^{C}=\operatorname{proj}(B_{i}^{C}\cap\mathbb{R}^{n}_{-})$, $\{\tilde{B}_{i}^{C}:\in J_{C}\}$ is a $\delta$-admissible family in $\mathbb{S}^{n-2}_{R_{C}}$ which is an $(n-2)$-sphere of radius $\sin(R_{C})$. By Proposition 4.9, if
@@
Define $K=K(n)=\frac{2}{\delta_{n-2}}$. Using that $\lim_{x\to 0}\frac{x}{\sin(x)}=1$, let $\eta_{0}>0$ be such that $0<x<\eta_{0}$ implies $20<\frac{x}{\sin(x)}<$ 2. Set $\overline{\delta}_{n}=\frac{\eta_{0}}{K}$. Hence if $0<\delta<\overline{\delta}_{n}$, $0<K\delta<\eta_{0}$ and by (11)
@@
  1. (E1)
    $F^{\prime}|_{E}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}^{*}}$.
  2. (E2)
    $F^{\prime}(x)\llcorner(\mathbb{D}^{n}\setminus B_{r_{E}}(e))=F(x)\llcorner(\mathbb{D}^{n}\setminus B_{r_{E}}(e))$ for every $x\in E$.
  3. (E3)
@@
  4. (E4)
    $U_{i}^{E}\cap T_{C}=\emptyset$ for every $i\in J_{E}$ and every $C\in\operatorname{Faces}_{p}(X)$ containing $E$.
  5. (E5)
@@
Base case. Let $k=1$ and let $E$ be a $1$-face of $X$. Same as for $n=3$ we set
$r_{E}=\max\{R_{C}:C\in\operatorname{Faces}_{p}(X),C\supseteq E\}$
being $L+K\delta\leq r_{E}\leq L+(K+1)\delta$ and $\partial B_{r_{E}}(e)\cap U^{E}_{i}=\emptyset$ for all $i\in I_{E}$. Define $J_{E}$ and $\{U^{E}_{i}:i\in J_{E}\}$ as in the inductive property. Let $x$ and $y$ be the vertices of $E$. Identify $E$ with the interval $[0,1]$ where $x\cong 0$ and $y\cong 1$ and set
$F^{\prime}(t)=F(t)\llcorner(\mathbb{D}^{n}\setminus B_{r_{E}}(e))+F_{E}(t)$
@@
Inductive step. We proceed as for $n=3$. The main difference is how to homotop $F_{E}^{0}:\partial E\to\mathcal{Z}_{0}(B_{E})$ defined as $F_{E}^{0}(x)=F^{\prime}(x)\llcorner B_{E}$ to a family $F_{E}^{1}$ for which $F_{E}^{1}\llcorner\partial B$ is supported in a discrete set which is independent of $x\in\partial E$.
@@
$B_{E}^{*}=B_{E}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}T_{C}$
@@
$S_{E}^{*}=S_{E}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}S_{C}$
where $S_{E}=B_{E}\cap\mathbb{S}^{n-1}$. Define also
$\mathbb{S}_{E}=\mathbb{S}^{n-1}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}\Pi_{C}.$
Each connected component of $\mathbb{S}_{E}$ is contractible. Let us pick points $y_{1},...,y_{N}\in\mathbb{S}_{E}$ one for each connected component and define a homotopy $R_{E}:\mathbb{S}_{E}\times[0,1]\to\mathbb{S}_{E}$ such that $R_{E}(x,0)=x$ and $R_{E}(x,1)=y_{j}$ if $x$ belongs to the same connected component as $y_{j}$. This yields a homotopy $H_{E}:B_{E}^{*}\times[0,1]\to B_{E}^{*}$ which is the identity outside of $R(\frac{L}{2},\varepsilon)$ and inside it has the following expression in coordinates $(u,t,s)\in\mathbb{S}^{n-1}\times[-\cos(\frac{L}{2}),-\cos(r_{E})]\times[0,\varepsilon]$:
@@
where $h(t)$ is the linear function such that $h(-\cos(\frac{L}{2}))=0$ and $h(-\cos(L))=1$. One can check that $H_{E}$ is continuous because for $s=\varepsilon$ and $t=-\cos(\frac{L}{2})$ it holds $H_{E}((u,t,s),t^{\prime})=(u,t,s)$. In addition, $\operatorname{Im}(H_{E})\subseteq B_{E}^{*}$ as the $u$-coordinate stays in $\mathbb{S}_{E}$, $H_{E}$ maps $\operatorname{int}(B_{E})\cup B$ into itself and same for $S_{E}^{*}\setminus B$. But moreover, we can see that
$H_{E}((S_{E}^{*}\setminus B)\times\{1\})=\{(y_{j},t,0):1\leq j\leq N,-\cos(L)\leq t\leq-\cos(r_{E})\}$
Let us define $F_{E}^{1}(x)$ to be the map obtained from $H_{E}(F_{E}^{0}(x),1)$ by applying the retraction $B_{E}^{*}\to\overline{B}$ given by $(u,t,s)\mapsto(u,t,g(s))$ with
$g(s)=\begin{cases}s&\text{if }s\leq-\cos(L)\\ -\cos(L)&\text{otherwise}.\end{cases}$
The underlying homotopy verifies properties (1) to (4) above with
for every $x\in\partial E$ (recall that $k=\dim(E)-1$ so this is stronger than (E5)). Additionally, $F_{E}^{1}(x)\llcorner\partial B$ is supported in $\{(y_{j},-\cos(L),0):1\leq j\leq N\}$ for every $x\in\partial E$.
Next we want to homotop $F_{E}^{1}$ into a family $F_{E}^{2}$ with the additional property that $F_{E}^{1}(x)\llcorner\partial B$ is supported in a fixed point for every $x\in\partial E$ (which will be chosen to be $\phi(y_{N},-\cos(L),0)$). We proceed as for $n=3$, defining $N-1$ homotopies $\overline{H}_{E}^{1},...,\overline{H}_{E}^{N-1}$ such that if $\overline{F}_{E}^{l}(x)$ is the map obtained after applying the first $l$ of them, then $\overline{F}_{E}^{l}(x)\llcorner\partial B$ is supported in $\{(y_{j},-\cos(L),0):j\geq l+1\}$.
We construct each $\overline{H}_{E}^{j}$ as follows. Denote $\overline{y}_{j}=(y_{j},-\cos(L),0)$. We consider an embedded curve $\gamma$ in $B_{E}^{*}$ connecting $\overline{y}_{j}$ with $\overline{y}_{j+1}$ and take a small tubular neighborhood $N_{\gamma}$ of $\gamma$ in $B_{E}^{*}$ which does not contain any $\overline{y}_{l}$ for $j\neq l,j+1$. Consider a diffeomorphism between $N_{\gamma}$ and $\mathbb{D}^{n-1}\times[0,1]$ and choose coordinates $(u,s)\in\mathbb{D}^{n-1}\times[0,1]$. We define $\overline{H}_{E}^{j}$ in $N_{\gamma}$ in a way so that points with $|u|\leq\frac{1}{2}$ are retracted to $\mathbb{D}^{n}\times\{1\}$ and in $\frac{1}{2}\leq|u|\leq 1$ we interpolate between that map and the identity as we approach $\partial\mathbb{D}^{n}\times[0,1]$ so that we can extend $\overline{H}_{E}^{j}$ by the identity in $B_{E}^{*}\setminus N_{\gamma}$. To be precise, we consider the following homotopy
@@
where $a:\mathbb{R}\to\mathbb{R}$ is a linear function with $a(\frac{1}{2})=1$ and $a(1)=0$. Then we can see that $\overline{H}_{E}^{j}$ is the identity in $\partial\mathbb{D}^{n-1}\times[0,1]\cup\mathbb{D}^{n-1}\times[1]$ providing a well defined homotopy in $B_{E}^{*}\cap B$. In addition, $\overline{H}_{E}^{j}((0,0),t)=(0,t)$ which means that the trajectory of $\overline{y}_{j}$ along $\overline{H}_{E}^{j}$ is precisely $\gamma$ as desired. Thus the number of points in $\operatorname{int}(\mathbb{D}^{n})\cup B$ gets increased by at most $1$ during the homotopy $\overline{H}_{E}^{j}(\overline{F}_{j-1}^{E}(x),t)$ and moreover at $t=1$ we have
@@
as the only point which may have entered that region during the homotopy is $\overline{y}_{j}$ which at the end of the homotopy overlaps with $\overline{y}_{j+1}$. We set $F_{E}^{2}=\overline{F}^{E}_{N-1}$, which by construction has the property that
@@
and verifies (1) to (4). If we apply a radial contraction of $B^{*}_{E}$ towards $e$ (specifically $(u,s,t)\mapsto(u,s,(1-t^{\prime})t+t^{\prime}(-1))$ for $t^{\prime}\in[0,1]$), we obtain a contraction of $F_{E}^{2}$ through families verifying (E1) to (E5), which allows to extend $F^{\prime}$ to all of $E$.
@@ 4.3. Parametric isoperimetric inequality in compact manifolds
Let $(M,g)$ be an $n$-dimensional compact triangulated Riemannian manifold with corners. Let $P\subseteq M$ be an $m$-dimensional subcomplex of $M$. Let $G:X^{p}\to\mathcal{I}_{1}(P,P\cap\partial M)$ be a continuous $\delta$-localized family of $1$-chains. Denote $F(x)=\partial G(x)$. Let $Q$ be a closed top dimensional cell of $P$ and let
@@
There exists a homotopy $G:X^{p}\times[0,1]\to\mathcal{I}_{1}(P,P\cap\partial M)$ with $\partial G(x,t)=F(x)$ for every $t\in[0,1]$ such that if $\overline{G}(x)=G(x,1)$ then
@@
Proof.
For simplicity, let us assume $\partial M=\emptyset$ (the case $\partial M\neq\emptyset$ can be handled by considering the double of $M$ as in Corollary 3.10). Let $m=\dim(P)$. We start by applying Lemma 3.9 to write $G(x)=G_{Q}(x)+G_{Q^{C}}(x)$ where both $G_{Q}$ and $G_{Q^{C}}$ are continuous families of absolute $1$-cycles and they are supported in $Q$ and $\operatorname{int}(Q)^{C}$ respectively. Let $F_{Q}(x)=\partial G_{Q}(x)$. In the following we identify $(Q,\partial Q)$ with $(\mathbb{D}^{m},\partial\mathbb{D}^{m})$ by a bilipschitz map with optimal Lipschitz constant. By Proposition 4.6, there exists a small ball $B\subseteq\partial Q$, a ball $\tilde{B}$ in $Q$ with $\tilde{B}\cap\partial Q\supseteq B$ and $\operatorname{rad}(\tilde{B})\leq\delta^{\prime}=c(n,p)\delta$ and a family $F_{Q}^{\prime}:X^{p}\to\mathcal{Z}_{0}(Q)$ which coincides with $F_{Q}$ outside of $\tilde{B}$ such that
@@
where $Q^{B}=\operatorname{int}(Q)\cup B$. As $F_{Q}^{\prime}-F_{Q}$ is supported in $\tilde{B}$ which is convex, by taking cones centered at a point in $\tilde{B}$ we can construct a family $R_{Q}:X^{p}\to\mathcal{I}_{1}(\tilde{B})$ such that $\partial R_{Q}(x)=F^{\prime}_{Q}(x)-F_{Q}(x)$. Let $S_{Q}(x)=G_{Q}(x)+R_{Q}(x)$ being $G_{Q}=S_{Q}+R_{Q}$. The advantage of this decomposition is that
@@
and hence $\operatorname{\mathbf{M}}(\partial S_{Q}(x)\llcorner Q^{B})\leq\operatorname{\mathbf{M}}_{Q}+p+1$. We will use this to construct a homotopy $S_{Q}:X^{p}\times[0,1]\to\mathcal{I}_{1}(Q)$ with $S_{Q}(x,0)=S_{Q}(x)$ such that
@@
    $\operatorname{\mathbf{M}}(S_{Q}(x,1)\llcorner\operatorname{int}(Q))\leq C(Q)(\operatorname{\mathbf{M}}_{Q}p^{-\frac{1}{n}}+p^{\frac{n-1}{n}})$, in fact $\overline{S}_{Q}(x)=S_{Q}(x,1)\llcorner\operatorname{int}(Q)$ is equal to the family constructed in Theorem 4.1 and hence it also has the property that $\operatorname{\mathbf{M}}(\partial\overline{S}_{Q}(x)\llcorner\partial Q)\leq 2(\operatorname{\mathbf{M}}_{Q}+p+1)$.
@@
We introduce the following notation. Given two points $z,w\in Q$ let $\Xi(z,w)$ be the segment connecting them. Given a 00-chain $x$ in $Q$ and a Lipschitz map $T:Q\to Q$ define
@@
where $\tilde{S}_{Q}(x)=S_{Q}(x,\frac{1}{2})$ and $\{\Psi_{t}:0\leq t\leq 1\}$ is an homotopy from $\Psi_{0}=id$ to the map $\Psi_{1}=\Psi$ from Lemma 4.4. It is clear that $\partial S_{Q}(x,t)=F_{Q}^{\prime}(x)$ for every $t\in[0,1]$. In addition, if $\overline{S}_{Q}(x)=S_{Q}(x,1)\llcorner\operatorname{int}(Q)$ as defined above then $\overline{S}_{Q}$ is equal to the map constructed in Theorem 4.1 with $r=p^{-\frac{1}{m}}$ and $F=F_{Q}^{\prime}$ (in particular, it is a continuous family of absolute $1$-chains), hence

### v2

@@ 4.2. Two important propositions to go from disks to manifolds
Given $C\in\operatorname{Faces}_{p}(X)$, we choose $r_{C}\in[L+K\delta,L+(K+1)\delta]$ ($K=K(n)$ to be determined later) such that $S^{n-1}_{r_{C}}\cap B_{i}^{C}=\emptyset$ for every $i\in I_{C}$ and $S^{n-1}_{r_{C}}\cap\mathcal{S}=\emptyset$ (being $\{B_{i}^{C}\}_{i\in I_{C}}$ the $\delta$-admissible family associated to $C$ and $\mathcal{S}=\bigcup_{x\in X_{0}}\operatorname{spt}(F^{\prime}(x))$). We denote
$J_{C}=\{i\in I_{C}:B^{C}_{i}\cap U_{r_{C}}=\emptyset\}.$
We claim that if we choose $K$ suitably, there exists a hyperplane $\Pi\subseteq\mathbb{R}^{n-1}$ such that $s_{\Pi}\cap B_{i}^{C}=\emptyset$ for every $i\in J_{C}$ and also $s_{\Pi}\cap\mathcal{S}=\emptyset$. Consider the map $\operatorname{proj}:A^{n-1}_{r_{C}}\to S^{n-2}_{r_{C}}$ given in $\phi$-coordinates by $\operatorname{proj}(u,t)=(u,-\cos(r_{C}))$. Notice that $\operatorname{proj}$ is $1$-Lipschitz and that $B_{i}^{C}\cap\mathbb{R}^{n}_{-}\subseteq S^{n-1}_{r_{C}}$ for every $i\in J_{C}$. Therefore, if $\tilde{B}_{i}^{C}=\operatorname{proj}(B_{i}^{C}\cap\mathbb{R}^{n}_{-})$, $\{\tilde{B}_{i}^{C}:\in J_{C}\}$ is a family of balls in $S^{n-2}_{r_{C}}$ with sum of radius less than $\delta$. As $S^{n-2}_{r_{C}}$ is an (n2)(n-2)-sphere of radius $\sin(r_{C})$, by Proposition 4.9, if
@@
Define $K=K(n)=\frac{2}{\delta_{n-2}}$. Using that $\lim_{x\to 0}\frac{x}{\sin(x)}=1$, let $\eta_{0}>0$ be such that $0<x<\eta_{0}$ implies $0<\frac{x}{\sin(x)}<2$. Set $\overline{\delta}_{n}=\frac{\eta_{0}}{K}$. Hence if $0<\delta<\overline{\delta}_{n}$, $0<K\delta<\eta_{0}$ and by (11)
@@
  1. (G1)
    $F^{\prime}|_{E}$ is localized in $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$.
  2. (G2)
    $F^{\prime}(x)\llcorner(\mathbb{D}^{n}\setminus U_{r_{E}})=F(x)\llcorner(\mathbb{D}^{n}\setminus U_{r_{E}})$ for every $x\in E$.
  3. (G3)
@@
  4. (G4)
    $B_{i}^{E}\cap T_{C}=\emptyset$ for every $i\in J_{E}$ and every $C\in\operatorname{Faces}_{p}(X)$ containing $E$.
  5. (G5)
@@
Base case. Let $k=1$ and let $E$ be a $1$-face of $X$. Same as for $n=3$, we set
$r_{E}=\max\{r_{C}:C\in\operatorname{Faces}_{p}(X),C\supseteq E\}.$
It follows $L+K\delta\leq r_{E}\leq L+(K+1)\delta$ and $S^{n-1}_{r_{E}}\cap B^{E}_{i}=\emptyset$ for all $i\in I_{E}$. Define $J_{E}$ and $\{B^{E}_{i}:i\in I_{E}^{*}\}$ as in the inductive property. Let $x$ and $y$ be the vertices of $E$. Identify $E$ with the interval $[0,1]$ where $x\cong 0$ and $y\cong 1$ and set
$F^{\prime}(t)=F(t)\llcorner(\mathbb{D}^{n}\setminus U_{r_{E}})+F_{E}(t)$
@@
Inductive step. Let $E$ be a $(k+1)$-face of $X$. Choose $r_{E}\in[L+(K+k)\delta,L+(K+k+1)\delta]$ so that $S^{n-1}_{r_{E}}\cap B_{i}^{E}=\emptyset$ for every $i\in I_{E}$ and $S^{n-1}_{r_{E}}\cap\mathcal{S}=\emptyset$. Define $J_{E}$ and $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$ as in the inductive property. Denote
$U_{E}$ $=U_{r_{E}},$
$B_{E}$ $=B_{r_{E}}.$
We proceed as for $n=3$. The main difference is how to homotop the continuous family $F_{E}^{0}:\partial E\to\mathcal{Z}_{0}(U_{E})$ defined as $F_{E}^{0}(x)=F^{\prime}(x)\llcorner U_{E}$ to a family $F_{E}^{1}$ for which $F_{E}^{1}\llcorner\partial B$ is supported in a discrete set which is independent of $x\in\partial E$ (recall that $B=B_{L}$ so $\partial B=S^{n-2}_{L}$).
@@
$U_{E}^{*}=U_{E}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}T_{C}$
@@
$B_{E}^{*}=B_{E}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}s_{C}$
(recall that $B_{E}=U_{E}\cap\mathbb{S}^{n-1}$). Define also
$\mathbb{S}^{n-2}_{E}=\mathbb{S}^{n-2}\setminus\bigcup_{\begin{subarray}{c}C\in\operatorname{Faces}_{p}(X)\\ C\supseteq E\end{subarray}}\Pi_{C}.$
Each connected component of $\mathbb{S}^{n-2}_{E}$ is contractible. Let us pick points $y_{1},...,y_{N}\in\mathbb{S}^{n-2}_{E}$ one for each connected component and define a Lipschitz homotopy $R_{E}:\mathbb{S}^{n-2}_{E}\times[0,1]\to\mathbb{S}_{E}$ such that $R_{E}(x,0)=x$ and $R_{E}(x,1)=y_{j}$ if $x$ belongs to the same connected component as $y_{j}$. This yields a homotopy $H_{E}:U_{E}^{*}\times[0,1]\to U_{E}^{*}$ which is the identity outside of $A^{n}_{\frac{L}{2},\varepsilon}$ and inside it has the following expression in coordinates $(u,t,s)\in\mathbb{S}^{n-1}\times[-\cos(\frac{L}{2}),-\cos(r_{E})]\times[0,\varepsilon]$:
@@
where $h(t)$ is the linear function such that $h(-\cos(\frac{L}{2}))=0$ and $h(-\cos(L))=1$. One can check that $H_{E}$ is continuous because for $s=\varepsilon$ and $t=-\cos(\frac{L}{2})$ it holds $H_{E}((u,t,s),t^{\prime})=(u,t,s)$. In addition, $\operatorname{Im}(H_{E})\subseteq U_{E}^{*}$ as the $u$-coordinate stays in $\mathbb{S}^{n-2}_{E}$, $H_{E}$ maps $(U_{E}\cap\operatorname{int}(\mathbb{D}^{n}))\cup B$ into itself and the same holds for $B_{E}^{*}\setminus B$. But moreover, we can see that
$H_{E}((B_{E}^{*}\setminus B)\times\{1\})=\{\phi(y_{j},t,0):1\leq j\leq N,-\cos(L)\leq t\leq-\cos(r_{E})\}$
Let us define $F_{E}^{1}(x)$ to be the map obtained from $H_{E}(F_{E}^{0}(x),1)$ by applying the retraction $U_{E}^{*}\to\overline{U}$ given in coordinates by $(u,t,s)\mapsto(u,g(t),s)$ with
$g(t)=\begin{cases}t&\text{if }t\leq-\cos(L)\\ -\cos(L)&\text{otherwise}.\end{cases}$
The underlying homotopy $F_{E}^{0}(x,t)$ from $F_{E}^{0}$ to $F_{E}^{1}$ verifies properties (G1) to (G4) above with
$\operatorname{\mathbf{M}}(F_{E}^{0}(x,t)\llcorner\operatorname{int}(\mathbb{D}^{n}))+\operatorname{\mathbf{M}}(F_{E}^{0}(x,t)\llcorner B)\leq M_{E}+k+1$
for every $x\in\partial E$ (recall that $k=\dim(E)-1$ so this is stronger than (G5)). Additionally, $F_{E}^{1}(x)\llcorner\partial B$ is supported in $\{\phi(y_{j},-\cos(L),0):1\leq j\leq N\}$ for every $x\in\partial E$.
Next we want to homotop $F_{E}^{1}$ into a family $F_{E}^{2}$ with the additional property that $F_{E}^{1}(x)\llcorner\partial B$ is supported in a fixed point for every $x\in\partial E$ (which will be chosen to be $\phi(y_{N},-\cos(L),0)$). We proceed as for $n=3$, defining N1N-1 homotopies $\overline{H}_{E}^{1},...,\overline{H}_{E}^{N-1}$ in $\overline{U}$ such that if $\overline{F}_{E}^{l}(x)$ is the map obtained after applying the first $l$ of them, then $\overline{F}_{E}^{l}(x)\llcorner\partial B$ is supported in $\{\phi(y_{j},-\cos(L),0):j\geq l+1\}$.
We construct each $\overline{H}_{E}^{j}$ as follows. Denote $\overline{y}_{j}=(y_{j},-\cos(L),0)$. We consider an embedded curve $\gamma$ in $S^{n-1}_{L}=\partial U\setminus B$ connecting $\overline{y}_{j}$ with $\overline{y}_{j+1}$ and take a small tubular neighborhood $N_{\gamma}$ of $\gamma$ in $\overline{U}$ which does not contain any $\overline{y}_{l}$ for $l\notin\{j,j+1\}$. Consider a diffeomorphism between $N_{\gamma}$ and $\mathbb{D}^{n-1}\times[0,1]$ and choose coordinates $(u,s)\in\mathbb{D}^{n-1}\times[0,1]$. We define $\overline{H}_{E}^{j}$ in $N_{\gamma}$ in a way so that points with $|u|\leq\frac{1}{2}$ are retracted to $\mathbb{D}^{n}\times\{1\}$ and in $\frac{1}{2}\leq|u|\leq 1$ we interpolate between that map and the identity as we approach $\partial\mathbb{D}^{n}\times[0,1]$ so that we can extend $\overline{H}_{E}^{j}$ by the identity in $\overline{U}\setminus N_{\gamma}$. To be precise, we consider the following homotopy
@@
where $a:\mathbb{R}\to\mathbb{R}$ is a linear function with $a(\frac{1}{2})=1$ and $a(1)=0$. Then we can see that $\overline{H}_{E}^{j}$ is the identity in $\partial\mathbb{D}^{n-1}\times[0,1]\cup\mathbb{D}^{n-1}\times[1]$ providing a well defined homotopy in $\overline{U}$. In addition, $\overline{H}_{E}^{j}((0,0),t)=(0,t)$ which means that the trajectory of $\overline{y}_{j}$ along $\overline{H}_{E}^{j}$ is precisely $\gamma$ as desired. Thus the number of points in $\operatorname{int}(\mathbb{D}^{n})\cup B$ gets increased by at most $1$ during the homotopy $\overline{H}_{E}^{j}(\overline{F}_{j-1}^{E}(x),t)$ and moreover at $t=1$ we have
@@
as the only point which may have entered that region during the homotopy is $\overline{y}_{j}$ which at the end of the homotopy overlaps with $\overline{y}_{j+1}$. We set $F_{E}^{2}=\overline{F}_{E}^{N-1}$, which by construction has the property that
@@
and verifies (G1) to (G4). If we apply a radial contraction of $\overline{U}$ towards $e$ (specifically $(u,s,t)\mapsto(u,s,(1-t^{\prime})t+t^{\prime}(-1))$ for $t^{\prime}\in[0,1]$), we obtain a contraction of $F_{E}^{2}$ through families verifying (G1) to (G5), which allows to extend $F^{\prime}$ to all of $E$.
@@ 4.3. Parametric isoperimetric inequality in compact manifolds
The goal of this section is to prove the Parametric Isoperimetric Inequality for families of 0-cycles supported in a Riemannian polyhedron (which yields the result in compact piecewise smooth Riemannian manifolds).
Theorem 4.18.
Let $P$ be an $m$-dimensional Riemannian polyhedron and let $P^{\prime}$ be a subpolyhedron. There exists a constant $C(P)$ depending only on $P$ such that the following holds. Let $F:X^{p}\to\mathcal{Z}_{0}(P,P^{\prime})$ be a contractible family of 0-cycles on $P$ and denote
Then there exists a continuous family $G:X^{p}\to\mathcal{I}_{1}(P,P^{\prime})$ such that
  1. (1)
    $\partial G(x)=F(x)$.
  2. (2)
We start with a contractible family $F:X^{p}\to\mathcal{Z}_{0}(P,\partial P^{\prime})$ supported in a Riemannian polyhedron $P$. We fix a smooth triangulation of $P$. In order to obtain a short filling $G$ of $F$, we will first start with a possibly very long filling $G:X^{p}\to\mathcal{Z}_{1}(M,\partial M)$, whose existence is guaranteed by the fact that $F$ is contractible. Then we will proceed to deform $G$ cell by cell, so that for each $m$-dimensional cell $Q$ of $M$ we homotop $G|_{Q}$ to the short filling $G_{Q}$ constructed in Theorem 4.1 (here we are using that $(Q,g)$ is bilipschitz homeomorphic to the Euclidean disk $\mathbb{D}^{n}$). The next proposition shows how to do such deformation. After doing that on each top dimensional cell $Q$ of $M$, $G$ could still have very large mass due to its part supported in the codimension-$1$ skeleton $P_{m-1}$ of $M$. Then what we do is to apply the same type of deformation at each (m1)(m-1)-cell of $M_{m-1}$, reducing the problem to the (m2)(m-2) skeleton of $M$. And then we proceed inductively. This is explained after proving 4.19, completing the proof of Theorem 4.18.
@@
Let $(P,g)$ be an $m$-dimensional Riemannian polyhedron. Let $G:X^{p}\to\mathcal{I}_{1}(P)$ be a continuous $\delta$-localized family of $1$-chains. Denote $F(x)=\partial G(x)$. Let $Q$ be a closed top dimensional cell of $P$ and let
@@
There exists a homotopy $G:X^{p}\times[0,1]\to\mathcal{I}_{1}(P)$ with $\partial G(x,t)=F(x)$ for every $t\in[0,1]$ such that if $\overline{G}(x)=G(x,1)$ then
@@
Proof of Proposition 4.19.
Let $m=\dim(P)$. We start by applying Lemma 3.8 to write $G(x)=G_{Q}(x)+G_{Q^{C}}(x)$ where both $G_{Q}$ and $G_{Q^{C}}$ are continuous families of absolute $1$-cycles and they are supported in $Q$ and $\operatorname{int}(Q)^{C}$ respectively. Let $F_{Q}(x)=\partial G_{Q}(x)$. In the following we identify $(Q,\partial Q)$ with $(\mathbb{D}^{m},\partial\mathbb{D}^{m})$ by a bilipschitz map with optimal Lipschitz constant. By Proposition 4.10, there exists a small ball $B\subseteq\partial Q$, a ball $\tilde{B}$ in $Q$ with $\tilde{B}\cap\partial Q\supseteq B$ and $\operatorname{rad}(\tilde{B})\leq\delta^{\prime}=c(n,p)\delta$ and a family $F_{Q}^{\prime}:X^{p}\to\mathcal{Z}_{0}(Q)$ which coincides with $F_{Q}$ outside of $\tilde{B}$ such that
@@
[…]
@@
and hence $\operatorname{\mathbf{M}}(\partial S_{Q}(x)\llcorner Q^{B})\leq\operatorname{\mathbf{M}}_{Q}+p+1$. Our goal is to replace the family $S_{Q}$ by a short family (one whose mass is bounded by $C(Q)(\operatorname{\mathbf{M}}_{Q}p^{-\frac{1}{m}}+p^{\frac{m-1}{m}})$) with boundary $F^{\prime}_{Q}$ using the construction in Theorem 4.1. To be precise, we will construct a homotopy $S_{Q}:X^{p}\times[0,1]\to\mathcal{I}_{1}(Q)$ with $S_{Q}(x,0)=S_{Q}(x)$ such that
@@
    $\operatorname{\mathbf{M}}(S_{Q}(x,1)\llcorner\operatorname{int}(Q))\leq C(Q)(\operatorname{\mathbf{M}}_{Q}p^{-\frac{1}{n}}+p^{\frac{n-1}{n}})$, in fact $\overline{S}_{Q}(x)=S_{Q}(x,1)\llcorner\operatorname{int}(Q)$ is equal to the family constructed in Theorem 4.1 (in particular, it is continuous) and hence it also has the property that $\operatorname{\mathbf{M}}(\partial\overline{S}_{Q}(x)\llcorner\partial Q)\leq 2(\operatorname{\mathbf{M}}_{Q}+p+1)$.
@@
We introduce the following notation. Given two points $z,w\in Q$ let $\Xi(z,w)$ be the segment connecting them. Given a 0-chain $x$ in $Q$ and a Lipschitz map $T:Q\to Q$ define
@@
where $\tilde{S}_{Q}(x)=S_{Q}(x,\frac{1}{2})$ and $\{\Psi_{t}:0\leq t\leq 1\}$ is an homotopy from $\Psi_{0}=id$ to the PL map $\Psi_{1}=\Psi$ from Lemma 4.4. It is clear that $\partial S_{Q}(x,t)=F_{Q}^{\prime}(x)$ for every $t\in[0,1]$. In addition, if $\overline{S}_{Q}(x)=S_{Q}(x,1)\llcorner\operatorname{int}(Q)$ as defined above then $\overline{S}_{Q}$ is equal to the map $G$ constructed in Theorem 4.1 with $r=p^{-\frac{1}{m}}$ and $F=F_{Q}^{\prime}$ (in particular, it is a continuous family of absolute $1$-chains), hence